ا = (هر را = (0) - (a) = ) = (ه) و = (الا ( 3 / 2 2 و و 2 گیا - ۔ 41e) = 4(e) = 0 24+y = fext @ 4

use the convolution theorem to solve the I.V.P.
Answer #1

y ty= 3 fins 9(a)= y (6) = 0 transfrom both sides Talets Laplace L2 yll (1 9 + Lycurga 3. L2 flash 12 YN-14 (6) - 4(o) + 37(O) 1 y+241+ y = fun og(u) = Lagenya FIA), let Talats lablace transform of eqr o both sides 2 y(x)) + 2 L2 y cury & Lycup =y + 2ylty = fone, O Y (0) =y! (a) = Taking Caplace transfrom both sides Lqylcuis + a Layllary + Lycay = LL Hary 12 // Al-Sy(